# Enhancing Security in Multi-Robot Systems through Co-Observation Planning, Reachability Analysis, and Network Flow

### v1

@@ III-C2 Cross-trajectory edges
Figure 4: Figure 4a, Latest departure node $q^{d}_{b}$ found for $v_{r1}$ and earliest arrival node $q^{a}_{b}$ found for $v_{r3}$. Figure 4b, example of a full checkpoint graph where round vertices are checkpoint generated through the heuristic search and triangle vertices are added with the cross-trajectory edges. Virtual source $v^{+}$ and sink $v^{-}$ are added to be used later in planning problem.
We first find trees of feasible paths between each security checkpoint and all other trajectories using position information alone (ignoring, for the moment, any timing constraint). More precisely, through $*\texttt{RRT}^{\texttt{*}}$, we find all feasible, quasi-optimal paths between one security checkpoint location $q_{v_{p}}(t_{v_{p}})$ where $v_{p}\in V_{p}$ for sub-team $\mathcal{I}_{p}$, and all the waypoints $\{q_{r}(t_{{r}_{i}})\}$ on any reference trajectory for a different sub-team $\mathcal{I}_{r}$. Then, to prune these trees, we consider the time needed to physically travel from one trajectory to the other while meeting other robots at the two endpoints. This is done by calculating the minimal travel time $t_{\textrm{path}}=\texttt{Cost}(q)/v_{max}$ for a robot to traverse each path. Two types of connecting nodes can be found.
@@
are waypoints on $\{q_{r}\}$ where robots from sub-team $\mathcal{I}_{p}$, deviating at $v_{p}$, can meet with sub-team $\mathcal{I}_{r}$ at $(q_{r}(t_{r_{i}}),t_{r_{i}})$. For these, $*\texttt{RRT}^{\texttt{*}}$ must have found a path from $v_{p}$ to $q_{r}(t^{a}_{r})$ with $t_{p}+t_{\textrm{path}}<t_{{r}_{i}}$, and $\mathcal{E}(q_{r}(t_{{r}_{i}}),q_{p},t_{{r}_{i}},t_{p})\cap F=\emptyset$.
@@
are the waypoints on $\{q_{r}\}$ such that a robot from sub-team $\mathcal{I}_{r}$, deviating from $(q_{r}(t_{{r}_{i}}),t_{{r}_{i}})$, can meet with robots in sub-team $\mathcal{I}_{p}$ at $v_{p}$. For these nodes, $*\texttt{RRT}^{\texttt{*}}$ must find a path from $v_{p}$ to $q_{r}(t_{{r}_{i}})$ for $v_{p}$ if $t_{p}>t_{{r}_{i}}+t_{\textrm{path}}$ and $\mathcal{E}(q_{p},q_{r}(t_{{r}_{i}}),t_{p},t_{{r}_{i}})\cap F=\emptyset$.
@@
For each $v_{p}\in V_{p}$ found through Algorithm 1, all the latest departure and earliest arrival nodes are added as vertices $V_{q}=V_{p}\cup\{v_{ea},v_{ld}\}$, and the corresponding paths are added as cross-trajectory edges to $E_{q}$. Examples are shown in Figure 4a.
III-C3 In-trajectory edges
We arrange all $V_{q}=\cup_{p}\{v^{0}_{p},\dots,v^{T}_{p}\}$ found in Sections III-C1 and III-C2 in ascending order of $t_{v^{i}_{p}}$. We then add to $E_{q}$ the in-trajectory edges $\{(v^{i}_{p}\rightarrow v^{i+1}_{p})\}$ obtained by connecting all consecutive checkpoints $v^{i}_{p}\in V_{q}$ with the checkpoints $v^{i+1}_{p}$ that follow them in their original trajectories. Examples are shown in Figure 4b.
@@ III-D Co-observation planning problem
In this section, we formulate the cross-trajectory planning problem as a network multi-flow problem, and solve it using mixed-integer linear programs (MILP). We assume that $n_{p}$ robots are dedicated (one in each sub-team) to follow the reference trajectory (named reference robots). The goal is to plan the routes of the $n-n_{p}$ additional cross-trajectory robots dedicated to cross-trajectory co-observations, and potentially minimize the number of ross-trajectory robots needed.
@@
Note that we assign fixed roles to robots for convenience in explaining the multi-flow formulation. In practice, after a cross-trajectory robot joins a team, it is considered interchangable, and could switch roles with the reference robot of that trajectory.
To formulate the problem as a network multi-flow problem, we augment the checkpoint graph $G_{q}$ to a flow graph $G=(V,E)$. The vertices of the new graph $G$ are defined as $V=V_{q}\cup\{v^{+},v^{-}\}$, where $v^{+}$ is a virtual source node, and $v^{-}$ is a virtual sink node. We add directed edges from $v^{+}$ to all the start vertices, and from all end vertices to $v^{-}$, with $v^{0}_{p}$ and $v^{T}_{p}$ representing the start and end vertices of sub-team $\mathcal{I}_{p}$. The edges of the new graph are defined as $E=E_{q}\cup\{(v_{+},v^{0}_{p})\}_{p}\cup\{(v^{T}_{p},v_{-})\}_{p}$.
The path of a robot $k$ all starts from $v^{+}$ and ends at $v^{-}$, and are represented as a flow vector $\mathbf{f}^{k}=\{f^{k}_{ij}\}$, where $f^{k}_{ij}\in\{1,0\}$ is an indicator variable representing whether robot $k$’s path contains the edge $v_{i}\to v_{j}$. The planning problem can be formulated as a path cover problem on $G_{q}$, i.e., as finding a set of paths $F=[\mathbf{f}_{1},\dots,\mathbf{f}_{\mathcal{K}}]$ for cross-trajectory robots such that every checkpoint in $\cup_{p}V_{p}$ is included in at least one path in $F$ (to ensure co-observation at every checkpoint as required by Remark 1).
Technically, we can always create a trivial schedule that involves only co-observations between members of the same team; this, however, would make the solution more vulnerable in the case where multiple agents are compromised in the same team. While in this paper we explicitly consider only the single attacker scenario, multi-attacks can be potentially handled by taking advantage of the decentralized blocklist protocol introduced in [3]. For this reason, we setup the methods presented below to always prefer cross-trajectory co-observation when feasible.
@@
With the formulation, the planning problem is written as an optimization problem, where the optimization cost balances between the co-observation performance and total number of flows (cross-trajectory robots) needed:
@@
The first term in the cost (55a) represents the total number of robot used, and the second term represents the overall co-observation performance (defined as the total number of cross-trajectory edges taken by all flows beyond the regular trajectory edges); the constant $\rho$ is a manually selected penalty parameter to balance between the two terms. (55b) is the flow conservation constraint to ensures that the amount of flow entering and leaving a given node $v$ is equal (except for $v^{+}$ and $v^{-}$). (55c) is the flow coverage constraints to ensures that all security checkpoints $\{V^{s}_{p}\}$ have been visited co-observed. The security graph $G_{q}$ is acyclic, making this problem in complexity class $P$, thus, can be solved in polynominal time [29].
@@ III-E Co-observation performance
Notice that problem (III-D) is guaranteed to have a solution for $\mathcal{K}=N_{p}$ where all $\mathcal{K}$ robots follows the reference trajectory ($f^{k}_{ij}=1,\forall\mathcal{I}_{v_{i}}=\mathcal{I}_{v_{j}}=k$). It is possible for the resulting flows to have a subset of flows $F_{e}\in F$ that is empty, i.e. $f_{ij}=0,\forall(v_{i},v_{j})\in E$, $f\in F_{e}$; these flow will not increase the cost and can be discarded from the solution.
The constant $\rho$ selects the trade-off between the number of surveillance robots and security performance. An increase in robots generally enhancing security via cross-trajectory co-observations; this also increases the complexity of the coordination across robots. In order to identify the minimum number of robots necessary, we propose an iterative approach where we start with $\mathcal{K}=1$ and gradually increases it until $F$ contains an empty flow, indicating the point where further robot additions do not improve performance.
@@
The value of row $\rho$ is upper bounded such that the second term for each single flow in (55a) is always smaller than one, i.e., $\rho\sum_{(ij)\in E}w_{ij}f^{k}_{ij}\leq 1$. Otherwise, the iteration continues indefinitely as adding additional flow introduce a negative term $\sum_{(+i)\in E}f^{k}_{+i}-\rho\sum_{(ij)\in E}w_{ij}f^{k}_{ij}=1-\rho\sum_{(ij)\in E}w_{ij}f^{k}_{ij}<0$ which always makes the cost (55a) smaller.
@@ III-F Result and simulation
Figure 5: Figure 5a showcase an unsecured 4-robot map exploration task. The cross-trajectory co-observation plan is shown as dotted arrors on top of the original plan in Figure 5a. The checkpoint graph and resulting flows are shown in Figure 5b.
We first test the proposed method for the example application in Figures 3 and 3c, using the same setup in Section II-I. Using the parameters $w_{c}=10$, $w_{t}=1$ and $\rho=0.01$, the result returns a total of $\mathcal{K}=3$ surveillance robots with cross-trajectory plan shown in Figure 5. The flows derived from the solution of the optimization problem (III-D) are highlighted in the graph Figure 5b, where each horizontal line represents the original trajectory of a sub-team and the number on each vertex $v_{i}$ represents the corresponding time $t_{i}$. The planning result in the workspace is shown in Figure 5a as dash-dotted arrows with the same color used for each flow in Figure 5b. Compared with the result in Figure 3c, it is easy to see that there is a significant improvement in map coverage in Figure 5a. At the same time, unsecured deviations to the forbidden regions of robots are secured through cross-trajectory observations.
The problem shows no solution for $\mathcal{K}\leq 2$. For cases $\mathcal{K}=3$, problem return the optimal result as shown in Figure 6. If we further increase $\mathcal{K}>3$, we do not get a better result; instead, the planner will return four flows with the rest $\mathcal{K}-3$ flows empty.
We then test the proposed method for a 4-team and 7-team case with result shown in Figure 6, where four trajectories are provided for a map exploration task with no security related constraints (co-observation schedule and reachability). We have a $10m\times 10m$ task space, three forbidden regions (rectangle regions in Figure 6a) and robots with a max velocity of $0.5m/dt$.
(a) 4-agent co-observation plan
(b) 7-agent co-observation plan
Figure 6: Cross-trajectory co-observation result of a 4 sub-teams task (Figure 6a) and a 7 sub-teams task Figure 6b
@@ IV Summary
This paper introduces security measures for protecting Multi-Robot Systems (MRS) from plan-deviation attacks. We establish a mutual observation schedule to ensure that any deviations into forbidden regions break the schedule, triggering detection. Co-observation requirements and reachability analysis are incorporated into the trajectory planning phase, formulated as constraints for the optimal problem. In scenarios where a secured plan is infeasible or enhanced task performance is necessary, we propose using redundant robots for cross-trajectory co-observations, offering the same security guarantees against plan-deviation attacks. However, the time-sensitive and proximity-dependent nature of co-observation necessitates a focus on collision avoidance during planning. Future work will integrate collision avoidance and explore dynamic duty assignments within sub-teams, increasing the challenge for potential attackers.
-A Proof of proposition 1
For subclaim 1a:
$H^{\mathrm{T}}H=H^{2}=4uu^{\mathrm{T}}uu^{\mathrm{T}}-4uu^{\mathrm{T}}+I^{2}=I,$ (56)
since $u^{\mathrm{T}}u=1$.
For subclaim 1b, let $U=\begin{bmatrix}u&u^{\bot}_{1},u^{\bot}_{2}\end{bmatrix}$, where $u^{\bot}_{1},u^{\bot}_{2}$ are two orthonormal vectors such that $I=UU^{\mathrm{T}}=uu^{\mathrm{T}}+u^{\bot}_{1}(u^{\bot}_{1})^{\mathrm{T}}+u^{\bot}_{2}(u^{\bot}_{2})^{\mathrm{T}}$; then, substituting $I$ in (4), we have that the eigenvalue decomposition of $H$ is given by
@@
For subclaim 2, first note that $Hu=2uu^{\mathrm{T}}u-u=u$. It follows that the sum of $\nu_{\mathcal{F}}$ and $\nu_{\mathcal{E}}$ is invariant under $H$:
$H(\nu_{\mathcal{F}}+\nu_{\mathcal{E}})=Hu\lVert\nu_{\mathcal{F}}+\nu_{\mathcal{E}}\rVert=u\lVert\nu_{\mathcal{F}}+\nu_{\mathcal{E}}\rVert=\nu_{\mathcal{F}}+\nu_{\mathcal{E}},$ (58)
@@
$H(\nu_{\mathcal{F}}-\nu_{\mathcal{E}})=2uu^{\mathrm{T}}(\nu_{\mathcal{F}}-\nu_{\mathcal{E}})-(\nu_{\mathcal{F}}-\nu_{\mathcal{E}})^{2}=-(\nu_{\mathcal{F}}-\nu_{\mathcal{E}}).$ (59)
@@
$H\nu_{\mathcal{F}}=\frac{1}{2}\bigl{(}H(\nu_{\mathcal{F}}+\nu_{\mathcal{E}})+H(\nu_{\mathcal{F}}-\nu_{\mathcal{E}})\bigr{)}=\nu_{\mathcal{E}}$ (60)
-B Proof or proposition 2
@@
$\begin{split}\dot{H}=&2HH^{\mathrm{T}}(\dot{u}u^{\mathrm{T}}+u\dot{u}^{\mathrm{T}})=2H(-\dot{u}u^{\mathrm{T}}+u\dot{u}^{\mathrm{T}})\\ =&-2H[[u]_{\times}\dot{u}]_{\times}\\ =&-2H\left[[u]_{\times}\frac{\left(I-uu^{\mathrm{T}}\right)\left(I-\nu_{\mathcal{F}}\nu_{\mathcal{F}}^{\mathrm{T}}\right)}{\left\|u^{\prime}\right\|\left\|\nu_{\mathcal{F}}\right\|}\dot{\nu}_{\mathcal{F}}\right]_{\times}\\ =&-2H[M\dot{\nu}_{\mathcal{F}}]_{\times}\!,\end{split}$ (63)
which is equivalent to the claim.
@@
$\begin{split}\dot{\pi}_{p\mathcal{E}}=&\dot{H}^{-1}SH(q_{avoid}-o)+H^{-1}\dot{S}H(q_{avoid}-o)\\ &+H^{-1}S\dot{H}(q_{avoid}-o)+(H^{-1}SH-I)\dot{o}\\ \end{split}$ (64)
@@
$H^{-1}\dot{S}H(q_{avoid}-o)=-H^{-1}Q^{\prime}q\dot{s}-sH^{-1}S^{2}\partial_{b}Qq\dot{b}\\ =\big{(}(q^{\mathrm{T}}\partial_{s}Q^{\prime}q)^{-1}H^{-1}Q^{\prime}qq^{\mathrm{T}}(4Q^{\prime}H[q_{avoid}-o]_{\times}U\\ +2Q^{\prime}H\partial_{q}o-\partial_{b}Q^{\prime}qq\partial_{q}b)-sH^{-1}S^{2}\partial_{b}Qq\partial_{q}b\big{)}\dot{q}\$ (74)
Thus equation (64) could be written as:

### v2

@@ III-C1 Cross-trajectory edges
Figure 6: (6a) The latest departure node $q^{d}_{b}$ found for $v_{r1}$ and the earliest arrival node $q^{a}_{b}$ found for $v_{r3}$. (6b) A full checkpoint graph, where circles are checkpoints $V_{p}$, triangles are connection nodes $V_{c}$, colored arrows (red and blue) are trajectory edges $E_{t}$ and black arrows are cross-trajectory edges $E_{c}$.
We first find trees of feasible paths between each security checkpoint and all other trajectories using position information alone (ignoring, for the moment, any timing constraint). More precisely, for each checkpoint $v_{p}\in\cup_{p}V_{p}$, we use $\texttt{RRT}^{\texttt{*}}$  to find all feasible, quasi-optimal paths from $q_{v_{p}}(t_{v_{p}})$ to all the waypoints $\{q_{r}(t_{{r}_{i}})\}$ on the reference trajectories of all other sub-teams $\mathcal{I}_{r}=\mathcal{I}\setminus\mathcal{I}_{p}$. Then, to prune these trees, we consider the time needed to physically travel from one trajectory to the other while meeting other robots at the two endpoints. This is done by calculating the minimal travel time $t_{\textrm{path}}=\texttt{Cost}(q)/v_{max}$ for a robot to traverse each path. Two types of connecting nodes can be found.
@@
are waypoints on $\{q_{r}\}$ where robots from sub-team $\mathcal{I}_{p}$, deviating at $v_{p}$, can meet with sub-team $\mathcal{I}_{r}$ at $(q_{r}(t_{r_{i}}),t_{r_{i}})$. For these, $\texttt{RRT}^{\texttt{*}}$ must have found a path from $v_{p}$ to $q_{r}(t^{a}_{r})$ with $t_{p}+t_{\textrm{path}}<t_{{r}_{i}}$, and $\mathcal{E}(q_{r}(t_{{r}_{i}}),q_{p},t_{{r}_{i}},t_{p})\cap F=\emptyset$.
@@
are the waypoints on $\{q_{r}\}$ such that a robot from sub-team $\mathcal{I}_{r}$, deviating from $(q_{r}(t_{{r}_{i}}),t_{{r}_{i}})$, can meet with robots in sub-team $\mathcal{I}_{p}$ at $v_{p}$. For these nodes, $\texttt{RRT}^{\texttt{*}}$ must find a path from $v_{p}$ to $q_{r}(t_{{r}_{i}})$ for $v_{p}$ if $t_{p}>t_{{r}_{i}}+t_{\textrm{path}}$ and $\mathcal{E}(q_{p},q_{r}(t_{{r}_{i}}),t_{p},t_{{r}_{i}})\cap F=\emptyset$.
@@
The set $V_{c}$ contains all the latest departure and earliest arrival nodes (if they are not identified as checkpoints already); the corresponding paths are added as cross-trajectory edges $E_{q}$. A toy example is shown in Fig. 6a.
III-C2 In-trajectory edges
We group $V=\{v^{i}_{p},\dots\}_{p=1}^{N_{p}}$ by sub-teams and arrange them in ascending order of their timestamps $t_{v^{i}_{p}}$. For each sub-team, consecutive vertices $v^{i}_{p}$ and $v^{i+1}_{p}$ are connected by adding in-trajectory edges $\{(v^{i}_{p}\rightarrow v^{i+1}_{p})\}$ to $E_{q}$, representing the corresponding segment of the planned trajectory. Examples are shown in Fig. 6b.
@@ III-D Co-observation planning problem
In this section, we formulate the cross-trajectory planning problem as a network multi-flow problem, and solve it using mixed-integer linear programs (MILP). We assume that $n_{p}$ robots are dedicated (one in each sub-team) to follow the reference trajectory (named reference robots). The goal is to plan the routes of the $n-n_{p}$ additional cross-trajectory robots dedicated to cross-trajectory co-observations, and potentially minimizing the number of cross-trajectory robots needed.
@@
Note that we assign fixed roles to robots for convenience in explaining the multi-flow formulation. In practice, after a cross-trajectory robot joins a team, it is considered interchangeable and could switch roles with the reference robot of that trajectory.
To formulate the problem as a network multi-flow problem, we augment the checkpoint graph to a flow graph. A virtual source node $v^{+}$ and a virtual source node are added to the vertices $V=(\cup_{p}V_{p})\cup V_{c}\cup\{v^{+},v^{-}\}$. Additional directed edges from $v^{+}$ to all the start vertices, and from all end vertices to $v^{-}$ are added $E=E_{q}\cup\{(v^{+},v^{0}_{p})\}_{p}\cup\{(v^{T}_{p},v^{-})\}_{p}$ with $v^{0}_{p}$ and $v^{T}_{p}$ representing the start and end vertices of sub-team $\mathcal{I}_{p}$ (dashed arrows in Fig. 1c).
The path of a robot $k$ all starts from $v^{+}$ and ends at $v^{-}$, and are represented as a flow vector $\mathbf{f}^{k}=\{f^{k}_{ij}\}$, where $f^{k}_{ij}\in\{1,0\}$ is an indicator variable representing whether robot $k$’s path contains the edge $v_{i}\to v_{j}$. The planning problem can be formulated as a vertex path cover problem on $G_{q}$, i.e., as finding a set of paths $F=[\mathbf{f}_{1},\dots,\mathbf{f}_{\mathcal{K}}]$ for cross-trajectory robots such that every checkpoint in $\cup_{p}V_{p}$ is included in at least one path in $F$ (to ensure co-observation at every checkpoint as required by Definition 4).
Technically, we can always create a trivial schedule that involves only co-observations between members of the same team; this, however, would make the solution more vulnerable in the case where multiple agents are compromised in the same team. In this paper we explicitly consider only the single-attacker scenario, multi-attacks can be potentially handled by taking advantage of the decentralized blocklist protocol introduced in [3]. For this reason, we setup the methods presented below to always prefer cross-trajectory co-observation when feasible.
@@
With the formulation, the planning problem is written as an optimization problem, where the optimization cost balances between the co-observation performance and the total number of flows (cross-trajectory robots) needed:
@@
The first term in the cost (38a) represents the total number of robots used, and the second term represents the overall co-observation performance (defined as the total number of cross-trajectory edges taken by all flows beyond the regular trajectory edges); the constant $\rho$ is a manually selected penalty parameter to balance between the two terms. The flow conservation constraint (38b) ensures that the amount of flow entering and leaving a given node $v$ is equal (except for $v^{+}$ and $v^{-}$). The flow coverage constraint (38c) ensures that all checkpoints $\{V^{s}_{p}\}$ have been visited and can support a co-observation. The checkpoint graph $G$ is acyclic, placing this problem in complexity class $P$; thus, it can be solved in polynomial time [33].
@@ III-E Co-observation performance
The network flow problem (38) is guaranteed to have a solution for $\mathcal{K}=N_{p}$ where all $\mathcal{K}$ robots follow the reference trajectory ($f^{k}_{ij}=1,\forall\mathcal{I}_{v_{i}}=\mathcal{I}_{v_{j}}=k$). It is possible for the resulting flows to have a subset of flows $F_{e}\in F$ that is empty, i.e. $f_{ij}=0,\forall(v_{i},v_{j})\in E$, $f\in F_{e}$; these flows will not increase the cost and can be discarded from the solution.
The constant $\rho$ selects the trade-off between the number of surveillance robots and security performance. An increase in robots generally enhances security via cross-trajectory co-observations; this also increases the complexity of the coordination across robots. To identify the minimum number of robots necessary, we propose an iterative approach in which we start with $\mathcal{K}=1$ and gradually increase it until $F$ contains an empty flow, indicating the point where further additions of robots do not improve performance.
@@
The value of row $\rho$ is upper-bounded such that the second term for each single flow in (38a) is always smaller than one, i.e., $\rho\sum_{(ij)\in E}w_{ij}f^{k}_{ij}\leq 1$. Otherwise, the iteration continues indefinitely as adding additional flow introduce a negative term $\sum_{(+i)\in E}f^{k}_{+i}-\rho\sum_{(ij)\in E}w_{ij}f^{k}_{ij}=1-\rho\sum_{(ij)\in E}w_{ij}f^{k}_{ij}<0$ which always makes the cost (38a) smaller.
@@ III-F Result and simulation
Figure 7: (7a) The cross-trajectory co-observation plan is shown as arrows on top of the original plan in Fig. 7a. (7b) The checkpoint graph and resulting flows highlighted.
We first test the proposed method for the example application in Figs. 4 and 4b, using the same setup in Section II-H. Using the parameters $w_{c}=10$, $w_{t}=1$ and $\rho=0.01$, the result returns a total of $\mathcal{K}=3$ surveillance robots with cross-trajectory plan shown in Fig. 7. The flows derived from the solution of the optimization problem (38) are highlighted in the graph Fig. 7b, where each horizontal line represents the original trajectory of a sub-team and the number on each vertex $v_{i}$ represents the corresponding time $t_{i}$. The planning result in the workspace is shown in Fig. 7a as dash-dotted arrows with the same color used for each flow in Fig. 7b. Compared with the result in Fig. 4b, it is easy to see that there is a significant improvement in map coverage in Fig. 7a. At the same time, unsecured deviations to the forbidden regions of robots are secured through cross-trajectory observations.
The problem shows no solution for $\mathcal{K}\leq 2$. For cases $\mathcal{K}=3$, the problem returns the optimal result as shown in Fig. 8. If we further increase $\mathcal{K}>3$, we do not get a better result; instead, the planner will return four flows with the rest $\mathcal{K}-3$ flows empty.
We then test the proposed method for 4-team and 7-team cases with results shown in Fig. 8, where four trajectories are provided for a map exploration task with no security-related constraints (co-observation schedule and reachability). We have a $10m\times 10m$ task space, three forbidden regions (rectangle regions in Fig. 8a), and robots with a max velocity of $0.5m/dt$.
(a) 4-agent case
(b) 7-agent case
Figure 8: (8a) Cross-trajectory co-observation result of a 4 sub-teams task. (8b) Cross-trajectory co-observation result of a 7 sub-teams task.
@@ IV Summary
This paper introduces security measures for protecting Multi-Robot Systems (MRS) from plan-deviation attacks. We establish a mutual observation schedule to ensure that any deviations into forbidden regions break the schedule, triggering detection. Co-observation requirements and reachability analysis are incorporated into the trajectory planning phase, and formulated as constraints for the optimal problem. In scenarios where a secured plan is infeasible or enhanced task performance is necessary, we propose using redundant robots for cross-trajectory co-observations, offering the same security guarantees against plan-deviation attacks. However, the time-sensitive and proximity-dependent nature of co-observation necessitates a focus on collision avoidance during planning. Future work will integrate collision avoidance and explore dynamic duty assignments within sub-teams, increasing the challenge for potential attackers.
-A Transformation of a reachability ellipsoid to canonical frame
The ellipse $\mathcal{E}$ expressed in $\mathcal{F}_{\mathcal{E}}$ is given by ${\mathcal{E}^{\mathcal{E}}}=\{{q^{\mathcal{E}}}\in m:d(q^{\mathcal{E}}_{1},{q^{\mathcal{E}}})+d({q^{\mathcal{E}}},q^{\mathcal{E}}_{2})<2a\}$, with foci $q^{\mathcal{E}}_{1},q^{\mathcal{E}}_{2}$ in $\mathcal{F}_{\mathcal{E}}$ defined as $q^{\mathcal{E}}_{1}=\begin{bmatrix}c&0&0\end{bmatrix}^{\mathrm{T}},q^{\mathcal{E}}_{2}=\begin{bmatrix}-c&0&0\end{bmatrix}^{\mathrm{T}}$, and semi-axis distance $c=\frac{\lVert q_{2}-q_{1}\rVert}{2}$.
Definition 8
The reachability ellipsoid $\mathcal{E}$ in the canonical frame is defined as the zero-level set of the quadratic function
where
$Q=\operatorname{diag}(a^{-2},b^{-2},b^{-2}),$ (40)
Lemma 1
${E^{\mathcal{F}}}(q)=(q-o_{\mathcal{E}})^{\mathrm{T}}H^{\mathrm{T}}QH(q-o_{\mathcal{E}})-1$ (41)
Proof:
-B Householder rotations
We define a differentiable transformation to a canonical ellipse that is used in the derivation of the reachability constraints in Sections II-C and II-G. This transformation includes a rotation derived from a modified version of Householder transformations [34]. We call our version of the operator a Householder rotation. In this section we derive Householder rotations and their differentials for the 3-D case; the 2-D case can be easily obtained by embedding it in the $z=0$ plane.
Definition 9
Let $\nu_{1}$ and $\nu_{2}$ be two unitary vectors ($\lVert\nu_{1}\rVert=\lVert\nu_{2}\rVert=1$). Define the normalized vector $u=\frac{\nu_{1}+\nu_{2}}{\lVert\nu_{1}+\nu_{2}\rVert}$, the Householder rotation $H(\nu_{1},\nu_{2})$ is defined as
$H(\nu_{1},\nu_{2})=2uu^{\mathrm{T}}-I.$ (42)
Here $H$ is a rotation mapping $\nu_{1}$ to $\nu_{2}$, as shown by the following.
Proposition 3
The matrix $H$ has the following properties:
  1. 1.
    It is a rotation, i.e. (a) $H^{\mathrm{T}}H=I$; (b) $\det(H)=1$
  2. 2.
    $\nu_{2}=H\nu_{1}$.
Proof:
For subclaim 1a, since $u^{\mathrm{T}}u=1$:
$H^{\mathrm{T}}H=H^{2}=4uu^{\mathrm{T}}uu^{\mathrm{T}}-4uu^{\mathrm{T}}+I^{2}=I.$ (43)
For subclaim 1b, let $U=\begin{bmatrix}u&u^{\bot}_{1}&u^{\bot}_{2}\end{bmatrix}$, where $u^{\bot}_{1},u^{\bot}_{2}$ are two orthonormal vectors such that $I=UU^{\mathrm{T}}=uu^{\mathrm{T}}+u^{\bot}_{1}(u^{\bot}_{1})^{\mathrm{T}}+u^{\bot}_{2}(u^{\bot}_{2})^{\mathrm{T}}$; then, substituting II in (42), we have that the eigenvalue decomposition of $H$ is given by
@@
For subclaim 2, first note that $Hu=2uu^{\mathrm{T}}u-u=u$. It follows that the sum of $\nu_{1}$ and $\nu_{2}$ is invariant under $H$:
$H(\nu_{1}+\nu_{2})=Hu\lVert\nu_{1}+\nu_{2}\rVert=u\lVert\nu_{1}+\nu_{2}\rVert=\nu_{1}+\nu_{2},$ (45)
@@
$H(\nu_{1}-\nu_{2})=2uu^{\mathrm{T}}(\nu_{1}-\nu_{2})-(\nu_{1}-\nu_{2})^{2}=-(\nu_{1}-\nu_{2}).$ (46)
@@
$H\nu_{1}=\frac{1}{2}\bigl{(}H(\nu_{1}+\nu_{2})+H(\nu_{1}-\nu_{2})\bigr{)}=\nu_{2}$ (47)
We compute the differential of $H$ implicitly using its definition (2). We use the notation $v]_{\times}\!:3\to 3\times 3$ to denote the matrix representation of the cross product with the vector $v$, i.e.,
such that $v]_{\times}w=v\times w$ for any $w\in 3$.
Proposition 4
Let $\nu_{1}(t)$ represent a parametric curve. Then
$\dot{H}=H[-2M\dot{\nu}_{\mathcal{F}}]_{\times}\!,$ (49)
where the matrix $M\in 3\times 3$ is given by
$M=[u]_{\times}\frac{\left(I-uu^{\mathrm{T}}\right)\left(I-\nu_{1}\nu_{1}^{\mathrm{T}}\right)}{\lVert u^{\prime}\rVert\lVert\nu_{1}\rVert}.$ (50)
Proof:
@@
$\begin{split}\dot{H}=&2HH^{\mathrm{T}}(\dot{u}u^{\mathrm{T}}+u\dot{u}^{\mathrm{T}})=2H(-\dot{u}u^{\mathrm{T}}+u\dot{u}^{\mathrm{T}})\\ =&-2H[[u]_{\times}\dot{u}]_{\times}\\ =&-2H\left[[u]_{\times}\frac{\left(I-uu^{\mathrm{T}}\right)\left(I-\nu_{1}\nu_{1}^{\mathrm{T}}\right)}{\left\|u^{\prime}\right\|\left\|\nu_{1}\right\|}\dot{\nu}_{\mathcal{F}}\right]_{\times}\\ =&-2H[M\dot{\nu}_{\mathcal{F}}]_{\times}\!,\end{split}$ (53)
which is equivalent to the claim. ∎
@@
[…]
@@
$H^{-1}\dot{S}H(q_{avoid}-o)=-H^{-1}Q^{\prime}q\dot{s}-sH^{-1}S^{2}\partial_{b}Qq\dot{b}\\ =\big{(}(q^{\mathrm{T}}\partial_{s}Q^{\prime}q)^{-1}H^{-1}Q^{\prime}qq^{\mathrm{T}}(4Q^{\prime}H[q_{avoid}-o]_{\times}U\\ +2Q^{\prime}H\partial_{q}o-\partial_{b}Q^{\prime}qq\partial_{q}b)-sH^{-1}S^{2}\partial_{b}Qq\partial_{q}b\big{)}\dot{q}\\$ (62)
Thus, equation (54) could be written as: